Mir Status Report 23 September 1997

Date: Tuesday, September 23, 1997 // MD 49/131

Mir Station System/Subsystem Activities -- The MCS is running normally and attitude control is being provided by 10 gyrodines. Elektron was reactivated today. Vozdukh repairs continued today, including installation of a newer fan and work on the cables. The crew reported at the DMT 04:15 that they had successfully activated the unit. No further work was performed on the urine water recovery system today.

Mir Payload Activities: Greenhouse Daily Ops (FE-1); Daily Exercise (Treadmill) with recording on 'ZY'; Greenhouse photography; MIM checkout; SAMS cleanup;


General Observations about the Seeds: All of the seeds that have germinated are showing cotyledon and are approximately a centimeter in height.

OPM was deactivated after the loss of MCS yesterday and was reactivated today at 15:00 DMT. MIM Checkout was completed successfully.

Camcorder - Mike confirmed the part number on the batteries is Sony Infolithium NP-F730. There is a charger currently on-orbit; three of these batteries will be manifested on the next Progress.

Stowage/Transfer - A radiogram with one change and two additions to the prepack list was submitted for uplink. Mike asked for clarification of the MIPS Harddrives to pack; he cannot find the serial numbers listed. Susan Minor relayed information back to the MOST to have Mike pack all of the hard drives he can find.

SMP/Medical - Mike participated in the final day of the physical training cycle with recording on the 'ZY' (monitored exercise).

Environmental Parameters (previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) BB 619 621 660-850
ppO2 (mmHg) 148 151 140-200
ppCO2 (mmHg) 4.7 5.7 up to 8.0
ppH20 (mmHg) 13.8 14.5 8-14
Humidity (percentage) 57 60 30-75
Temperature (xC) (BB) 24.6 25.8 18-28

  • submit to reddit


49th International Conference on Environmental Systems - ICES 2019

Join our mailing list

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on September 23, 1997 8:39 PM.

Mir Status Report 22 September 1997 was the previous entry in this blog.

Mir Status Report 24 September 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.